Find four consective odd integers whose sum is 112
Answer by Paul(988)

x+x+2+x+4+x+6=112
4x+12=112
4x=100
x=25
25+2=27, 25+4=29, 25+6=31
Hence, the 4 numbers are 25, 27, 29, and 31.